OpenStudy (anonymous):

how do you find the solutions to 1-tanX/2=0?

OpenStudy (anonymous):

start with \[\tan(\frac{x}{2})=1\] and then think what values give 1 as a tangent

OpenStudy (anonymous):

for example \[\tan(\frac{\pi}{4})=1\] so if \[2x=\frac{\pi}{4}\] then \[x=\frac{\pi}{8}\] that would be one answer. there are infinitely manyh

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

what about (1-tanx)/2=0